ADNAN MUHAMMED

Grade 12,

A car starting from rest accelerates at rate a through distance d,then continues at constant speed for time t,decelerates at the rate of a/2 to come to rest .if the total distance traveled is 15d,then d equals to

A car starting from rest accelerates at rate a through distance d,then continues at constant speed for time t,decelerates at the rate of a/2 to come to rest .if the total distance traveled is 15d,then d equals to 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear KArthik
 
 

S=(ft^2)/2

V.=squere root of(2Sf)

During retardation S2=2S

During constant velocity  15S -3S = 12S = vo

. '' .S= (ft^2)/72